# Predicting Time Pressure of Powered Two-Wheeler Riders for Proactive Safety Interventions

### v1

@@ Abstract
Time pressure critically influences risky maneuvers and crash proneness among powered two-wheeler riders, yet its prediction remains underexplored in intelligent transportation systems. We present a large-scale dataset of 129,000+ labeled multivariate time-series sequences from 153 rides by 51 participants under No, Low, and High Time Pressure conditions. Each sequence captures 63 features spanning vehicle kinematics, control inputs, behavioral violations, and environmental context. Our empirical analysis shows High Time Pressure induces 48% higher speeds, 36.4% greater speed variability, 58% more risky turns at intersections, 36% more sudden braking, and 50% higher rear brake forces versus No Time Pressure. To benchmark this dataset, we propose MotoTimePressure, a deep learning model combining convolutional preprocessing, dual-stage temporal attention, and Squeeze-and-Excitation feature recalibration, achieving 91.53% accuracy and 98.93% ROC AUC, outperforming eight baselines. Since time pressure cannot be directly measured in real time, we demonstrate its utility in collision prediction and threshold determination. Using MTPS-predicted time pressure as features, improves Informer-based collision risk accuracy from 91.25% to 93.51%, approaching oracle performance (93.72%). Thresholded time pressure states capture rider cognitive stress and enable proactive ITS interventions, including adaptive alerts, haptic feedback, V2I signaling, and speed guidance, supporting safer two-wheeler mobility under the Safe System Approach.
@@ I Introduction
Powered two-wheelers (PTWs) are among the most accessible and widely used transport modes, particularly in developing countries. However, they carry elevated crash risks due to exposure to rider and balance control demands [10122165]. The severity of the crash is strongly influenced by overspeeding, traffic dynamics, rider behavior, and road conditions [Seefong2024, 7328721, 9625971], with human error and risky maneuvers being the leading causes [pavlidis2016dissecting, 6899632, CHOUHAN20241378]. PTWs account for 21% of global road fatalities rising to 38% of deaths and 51% of injuries in Low- and Middle-Income Countries (LMICs) [WHO2023]. Almost half of traffic deaths involve vulnerable users such as pedestrians, cyclists, and PTW riders [WHO2023]. Excess speed alone contributes to 34% PTW fatalities, with riders traveling up to 2.3 times faster than car drivers [Sharma21042025].
@@
According to India’s Ministry of Road Transport and Highways (MoRTH), PTWs are the most prevalent and economical mode of transport in LMICs. Between 2003 and 2022, India’s registered vehicle population grew to 354 million, of which PTWs constituted 263 million (74.4%). This segment grew at a CAGR of 8.6%, surpassing buses (2.48%) and goods vehicles (7.3%). Category-wise registration trends are shown in Fig. 1, underscoring the growing dependence on PTWs and the urgent need for targeted safety interventions [MORTH202425].
Figure 1: Registered motor vehicles in India by category (2003–2022) [MORTH202425].
@@
PTWs also account for the largest share of traffic fatalities in India, accounting for 44.5% of deaths in 2022, followed by pedestrians (19.5%), cars/LMVs (12.5%), trucks/lorries (6.3%), autorickshaws (3.9%), bicycles (2.9%), buses (2.4%), and others (8.0%) [MORTH202324, MORTH202122, MORTH202223]. According to official MoRTH reports on road accidents in india , road fatalities in India are predominantly male, ranging from 85.2% to 87.3% between 2019 and 2023 [morth2019, morth2020, morth2021, morth2022, morth2023], with most fatalities occurring in the 18–45 age group. These statistics show the high risk associated with male PTW riding behaviour. Driving errors are the main cause of crash globally, especially for PTWs, which face high instability and dynamic control challenges [8294048, BHALLA201583]. Field studies of PTW behavior have shown significant variability in acceleration and braking patterns [Mondal21042023], emphasizing the inherent dynamic control challenges these vehicles face in real-world traffic. Simulator studies further reveal context-dependent rider behavior in mixed traffic [10073653], underscoring PTWs vulnerability and the need for targeted safety interventions.
@@
A critical yet underexplored factor in PTW safety is time pressure (TP), reflecting perceived urgency or limited time to complete riding tasks [PAWAR2022105582]. For example, many delivery PTW riders in India are subjected to TPs [fortune2025quickcommerce, thehindu2022delivery, jakartapost2022delivery]. Elevated TP impairs cognition and motor control, driving overspeeding, braking inconsistencies, delayed hazard response, and higher crash risk [PAWAR2022105582, Gupta25112024, Sharma21042025]. Naturalistic studies link risky behavior to speed, maneuvering, and context [KONG2020105620], while stress further amplifies variability and control errors [pavlidis2016dissecting]. Field evidence from Thailand also confirms that widespread speeding significantly raises collision risk [Seefong2024].
@@
Pawar et al. [PAWAR2022105582] showed that TP induces risky maneuvers. However, directly measuring TP in real-world PTW riding is practically impossible. We hypothesize that TP can be predicted from PTW’s data. Therefore, we established a proxy measure for TP and predict it using other PTW telematic features such as speed, acceleration, etc. Towards this end, we first collect data in the same fashion as [PAWAR2022105582] but for PTW. That is, riders experienced TP conditions (no time pressure (NTP), low time pressure (LTP), high time pressure (HTP)) and asked to drive PTW. Thus, we collect the supervised data. To the best of our knowledge, this is the first work that develops predictive models for cognitive state detection in PTW riders, addressing a fundamental limitation in current ITS. To evaluate the data quality, we propose a novel MotoTimePressure (MTPS) model that predicts TP. To show the utility of such a predicted TP feature in collision risk prediction, we train an Informer [zhou2021informer] model using the ground truth (GT) TP feature and the predicted TP feature. Our results show that predicted TP features increase the accuracy of collision risk prediction from 91.25% to 93.51%, approaching oracle performance (93.72%). The key contributions of this work are summarized as follows:
@@
    The First Labeled Dataset for PTW Rider Cognitive State: To the best of our knowledge, We present a first high-resolution PTW simulator dataset comprising more than 129,000 labeled multivariate time-series sequences from 153 rides by 51 participants under three TP conditions (NTP, LTP, HTP). Each sequence includes 63 features covering vehicle kinematics, control inputs, behavior violations, and environmental context, providing a rich resource for modeling cognitive load and rider safety.
@@
    Speed and Control Degradation Under TP: We provide comprehensive statistical evidence that TP systematically degrades riding control. Under HTP conditions, riders exhibit 48% higher mean speeds and 36.4% greater speed variability compared to NTP. Control degradation extends to 604% increased front brake force application, 50% higher rear brake forces, and 35% increased gear usage, demonstrating systemic impairment of fundamental riding skills under TP.
@@
    Systemic Risk Escalation and Inadequate Compensation: We demonstrate that TP creates compound collision risks across multiple riding domains. HTP causes: 58% more risky turns at incorrect speeds, 36% more sudden braking incidents, 50% higher brake forces, and 67% worse clutch control. Despite natural compensation behaviors, human adaptation fails as HTP still causes 21.3% worse control variability and 10.2% more severe safety violations. Statistical analysis reveals a consistent degradation pattern: TP $\rightarrow$ degraded control $\rightarrow$ insufficient self-correction $\rightarrow$ persistent collision risks.
@@
    A Novel Predictive Model for Time Pressure: We propose MTPS, the first architecture specifically designed for rider TP state prediction. MTPS introduces a novel integration of convolutional preprocessing, dual temporal attention blocks, and hierarchical feature fusion to capture multi-scale dependencies in riding behavior.
  5. 5.
    A New Paradigm for Proactive Safety: MTPS outperforms eight ML and DL baselines for TP classification in PTWs, achieving 91.53% accuracy and 98.93% ROC AUC. Incorporating MTPS-predicted TP into an Informer-based crash risk predictor increases the accuracy from 91.25% (baseline) to 93.51%, approaching oracle performance (93.72%) with GT TP.
@@ II-A Time Pressure and Driving Behaviour
Driving under TP can reduce the driver’s focus, which can lead to irrational decisions and a higher risk of collision [PAWAR2022105582, Gupta25112024]. Drivers operating under TP exhibit higher speeds, altered driving behaviour, shorter accepted gaps, increased cognitive workload, faster decision cycles, and higher crash risk [PAWAR20221, PAWAR2022105582, Gupta25112024, Sharma21042025, fortune2025quickcommerce, thehindu2022delivery, jakartapost2022delivery]. Nishant et al. [PAWAR202229] reported that TP significantly increases longitudinal acceleration, braking force variability, and steering fluctuations in a fixed-base simulator, reinforcing that TP elicits statistically significant shifts toward risk-oriented control behaviour.
TP is fundamentally a perception of urgency or limited time to complete the driving task. This perceived urgency narrows the attentional bandwidth and accelerates cognitive processing, often forcing drivers and riders to make rapid decisions with reduced situational awareness [Gupta25112024, Sharma21042025, PAWAR202229]. Although early work primarily focused on four-wheelers [pavlidis2016dissecting, Gupta25112024], recent PTW research reveals similar but more sensitive behavioural deterioration under TP, including inconsistent braking, excessive speeding, unstable motor control, and delayed hazard response [Sharma21042025, PAWAR2022105582]. These PTW-specific vulnerabilities are important because PTWs exhibit inherently higher crash risk and lower protection levels. Simulator-based studies further show that TP increases crash likelihood across multiple riding tasks. In complex maneuvers such as intersections, Pawar & Velaga [PAWAR2022105582] reported that TP increases cognitive urgency, speed and reduces caution. By narrowing attention and biasing riders toward fast, risk-oriented strategies, TP directly reduces safety margins and increases the probability of unsafe decisions.
@@ II-B Time Pressure and Gap-Acceptance Behaviour
Research on un-signalized intersections, particularly in LMIC, mainly focused on gap-acceptance behavior and critical gap estimation [cjce-2018-0796, 2461-16, PAWAR2022105582]. Gap acceptance is highly sensitive to TP [PAWAR2022105582]. Pawar et al. [PAWAR2022105582] showed that drivers under LTP and HTP accept significantly smaller gaps, resulting in elevated crash likelihoods of 127% (LTP) and 181% (HTP) relative to NTP. Their Generalized Linear Mixed Models (GLMM) and decision-tree (DT) analyses revealed that the TP effects on gap selection are jointly shaped by merge distance, oncoming vehicle speed, speed reduction before merging, and driver experience. These results illustrate how TP systematically modifies perceptual judgment and risk thresholds, thereby reshaping tactical decision-making at intersections.
@@ II-C Time Pressure Effects on PTW Riding: Speed and Over-Speeding
PTWs are particularly vulnerable in traffic, with over-speeding recognized as a major cause of PTW-related crashes [Sharma21042025, GUPTA2022105820, KONG2020105620, PAWAR2022105582]. Steg and van Brussel [STEG2009503] found that subjective norms and favourable attitudes toward speeding significantly increased young moped riders to over-speed, illustrating the role of cognitive and social influences.
Other recent PTW-specific studies show that TP amplifies these vulnerabilities. TP increases over-speeding tendencies due to psychological constructs such as travel-time anxiety, hurriedness, and perceived pressure to reach early [Seefong2024, Balasubramanian01022021, ATOMBO2016104, Sharma21042025, GUPTA2022105820, Michaels2014]. Latent class analyses identify three psychologically distinct rider segments minimally anxious, moderately anxious, and highly anxious, each showing different sensitivity to TP, with highly anxious riders most likely to over-speed. These findings highlight the need to integrate psychological and behavioural information when modelling TP-related riding risk.
@@ II-D Time Pressure and Distraction in MTW Riders
TP represents a cognitive state in which the perceived time available is insufficient to complete the riding task [GUPTA2022105820, RENDONVELEZ2016150, KONG2020105620, PAWAR2022105582]. Under TP, riders must process information more rapidly, often with reduced cognitive reserves. This accelerated decision environment narrows attention, degrades situational awareness, and increases susceptibility to operational errors [Leung01112012, PAWAR2020105401]. Although distraction and TP are distinct constructs, their interaction helps reveal how cognitive stress reshapes rider attention. Gupta et al. [GUPTA2022105820] found that TP can suppress intentional distractions (e.g., voluntary mobile use) due to heightened task focus, while simultaneously amplifying unintentional cognitive lapses driven by stress and mental workload. These findings reinforce that TP fundamentally alters cognitive processing, attention allocation, and behavioural stability factors central to TP prediction.
@@ II-E Simulator Validity Under Time Pressure
Driving simulators provide a safe, controlled environment for studying TP effects, eliminating real-world collision risks while maintaining experimental validity [PAWAR202229, Bham04032018, LI2019288]. Research consistently demonstrates that simulator behavior under TP closely mirrors real-world responses exhibiting what is termed relative validity [PAWAR202229]. These studies show that although absolute behavioral magnitudes may differ between simulated and natural environments, the fundamental patterns of TP-induced changes remain consistent: increased speeds, elevated braking intensity, and amplified vehicular variability occur in both settings. This evidence confirms that simulators are well-suited for TP research when predictive models account for systematic calibration differences between simulated and real-world data.
@@ II-F Machine Learning for Behaviour and Risk Modeling
In parallel to behavioural research, the ITS community has increasingly adopted ML and DL approaches to model driving risk. Traditional classifiers such as Random Forests (RF), Support Vector Machines (SVM), and adaptive Deep Neural Networks (DNNs) have been used to predict high-risk manoeuvres and collision likelihood, often reporting accuracies in the range of 84–90% [Bouhsissin2023, garefalakis2024predicting, Rodegast2024, Aci2025]. More advanced sequence models, including the TST [10.1609/aaai.v37i9.26317], Informer architecture [zhou2021informer], have demonstrated a strong capability to handle long time-series dependencies relevant to driving dynamics. Despite their predictive success, these models primarily estimate external outcomes (e.g., collisions, hazardous events) and remain largely insensitive to the internal cognitive precursors such as TP, that trigger risky behaviours. This highlights an important methodological gap: existing ML systems detect risk after behavioural deviations occur rather than predicting the underlying cognitive state that gives rise to those deviations.

### v3

@@ Abstract
Time pressure critically influences risky maneuvers and crash proneness among powered two-wheeler riders, yet its prediction remains underexplored in intelligent transportation systems. We present a large-scale dataset of 129,000+ labeled multivariate time-series sequences from 153 rides by 51 participants under No, Low, and High Time Pressure conditions. Each sequence captures 63 features spanning vehicle kinematics, control inputs, behavioral violations, and environmental context. Our empirical analysis shows High Time Pressure induces 48% higher speeds, 36.4% greater speed variability, 58% more risky turns at intersections, 36% more sudden braking, and 50% higher rear brake forces versus No Time Pressure. To benchmark this dataset, we propose MotoTimePressure, a deep learning model combining convolutional preprocessing, dual-stage temporal attention, and Squeeze-and-Excitation feature recalibration, achieving 91.53% accuracy and 98.93% ROC AUC, outperforming six baselines, with only 172K parameters, 0.66 MB model size, and 0.21 ms inference on CPU. Since time pressure cannot be directly measured in real time, we demonstrate its utility in collision prediction and threshold determination. Using MTPS-predicted time pressure as a feature improves collision risk accuracy for both Informer (91.25% to 93.51%) and TimesNet (92.10% to 93.90%), approaching oracle performance (93.72% and 94.06%, respectively). Thresholded time pressure states capture rider cognitive stress and enable proactive ITS interventions, including adaptive alerts, haptic feedback, V2I signaling, and speed guidance, supporting safer two-wheeler mobility under the Safe System Approach.
@@ I Introduction
Powered two-wheelers (PTWs) are among the most accessible and widely used transport modes, particularly in developing countries. However, they carry elevated crash risks due to exposure to rider and balance control demands [5]. The severity of the crash is strongly influenced by overspeeding, traffic dynamics, rider behavior, and road conditions [38, 4, 8], with human error and risky maneuvers being the leading causes [28, 2, 9]. PTWs account for 21% of global road fatalities rising to 38% of deaths and 51% of injuries in Low- and Middle-Income Countries (LMICs) [44]. Almost half of traffic deaths involve vulnerable users such as pedestrians, cyclists, and PTW riders [44]. Excess speed alone contributes to 34% PTW fatalities, with riders traveling up to 2.3 times faster than car drivers [39].
@@
According to India’s Ministry of Road Transport and Highways (MoRTH), PTWs are the most prevalent and economical mode of transport in LMICs. Between 2003 and 2022, India’s registered vehicle population grew to 354 million, of which PTWs constituted 263 million (74.4%). This segment grew at a CAGR of 8.6%, surpassing buses (2.48%) and goods vehicles (7.3%). Category-wise registration trends are shown in Fig. 1, underscoring the growing dependence on PTWs and the urgent need for targeted safety interventions [26].
Figure 1: Registered motor vehicles in India by category (2003–2022) [26].
@@
PTWs also account for the largest share of traffic fatalities in India, accounting for 44.5% of deaths in 2022, followed by pedestrians (19.5%), cars/LMVs (12.5%), trucks/lorries (6.3%), autorickshaws (3.9%), bicycles (2.9%), buses (2.4%), and others (8.0%) [18, 16, 17]. According to official MoRTH reports on road accidents in india, road fatalities in India are predominantly male, ranging from 85.2% to 87.3% between 2019 and 2023 [21, 22, 23, 24, 25], with most fatalities occurring in the 18–45 age group. These statistics show the high risk associated with male PTW riding behaviour. Driving errors are the main cause of crash globally, especially for PTWs, which face high instability and dynamic control challenges [3, 6]. Field studies of PTW behavior have shown significant variability in acceleration and braking patterns [20], emphasizing the inherent dynamic control challenges these vehicles face in real-world traffic. Simulator studies further reveal context-dependent rider behavior in mixed traffic [48], underscoring PTWs vulnerability and the need for targeted safety interventions.
@@
A critical yet underexplored factor in PTW safety is time pressure (TP), reflecting perceived urgency or limited time to complete riding tasks [32]. For example, many delivery PTW riders in India are subjected to TPs [40, 35, 36]. Elevated TP impairs cognition and motor control, driving overspeeding, braking inconsistencies, delayed hazard response, and higher crash risk [32, 11, 39]. Naturalistic studies link risky behavior to speed, maneuvering, and context, while stress further amplifies variability and control errors [28]. Field evidence from Thailand also confirms that widespread speeding significantly raises collision risk [38]. However, existing studies have only examined TP as a cause of risky behavior, not as a predictable cognitive state that can be detected in real-time.
@@
Pawar et al. [32] showed that TP induces risky maneuvers, but directly measuring TP in real-world PTW riding remains practically impossible. We hypothesize that TP can be predicted from PTW’s data. Therefore, we established a proxy measure for TP and predict it using other PTW telematic features such as speed, acceleration, etc. Towards this end, we first collect data in the same fashion as [32] but for PTW. That is, riders experienced TP conditions (no time pressure (NTP), low time pressure (LTP), high time pressure (HTP)) and asked to drive PTW. Thus, we collect the supervised data. To the best of our knowledge, this is the first work that develops predictive models for cognitive state detection in PTW riders, addressing a fundamental limitation in current ITS. To evaluate the data quality, we propose a novel MotoTimePressure (MTPS) model that predicts TP. To show the utility of such a predicted TP feature in collision risk prediction, we train an Informer [49] model using the ground truth (GT) TP feature and the predicted TP feature. Our results show that predicted TP features increase the accuracy of collision risk prediction from 91.25% to 93.51% for Informer and from 92.10% to 93.90% for TimesNet, approaching oracle performance (93.72% and 94.06%, respectively). The key contributions of this work are summarized as follows:
@@
    The First Labeled Dataset for PTW Rider Cognitive State: To the best of our knowledge, we present the first high-resolution PTW simulator dataset comprising more than 129,000 labeled multivariate time-series sequences from 153 rides by 51 participants under three TP conditions (NTP, LTP, HTP).
@@
    Systematic Behavioral Degradation Under TP: TP systematically degrades riding control and creates compound collision risks. Under HTP, riders exhibit 48% higher speeds, 36.4% greater variability, 604% increased front brake, 50% higher rear brake, 35% increased gear usage, 58% more risky turns, 36% more sudden braking, and 67% worse clutch control. Despite natural compensation, human adaptation fails — HTP still causes 21.3% worse control variability and 10.2% more violations, following: TP $\rightarrow$ degraded control $\rightarrow$ insufficient correction $\rightarrow$ persistent risk.
@@
    Predictive Model for Time Pressure: We propose MTPS, the first architecture specifically designed for rider TP state prediction. MTPS introduces a novel integration of convolutional preprocessing, dual temporal attention blocks, and hierarchical feature fusion to capture multi-scale dependencies in riding behavior, achieving 91.53% accuracy and 98.93% ROC AUC, outperforming six ML and DL baselines including TST, SHGN, and Informer. Moreover, MTPS achieves 0.21 ms inference on CPU with only 172K parameters and 0.66 MB size, enabling real-time proactive safety on edge devices.
@@
    A New Paradigm for Proactive Safety: Incorporating MTPS-predicted TP into crash risk predictors improves accuracy from 91.25% to 93.51% for Informer and from 92.10% to 93.90% for TimesNet, approaching oracle performance (93.72% and 94.06%, respectively). This demonstrates that TP provides cognitive state information beyond kinematics, enabling pre-crash intervention.
@@ II Literature Review
To establish the foundation for our work, this section reviews existing research on TP in driving and riding contexts, gap-acceptance behavior, PTW-specific vulnerabilities, simulator validity, and machine learning approaches to risk modeling.
@@ II-A Time Pressure and Driving Behaviour
Driving under TP can reduce the driver’s focus, which can lead to irrational decisions and a higher risk of collision [32, 11]. Drivers operating under TP exhibit higher speeds, altered driving behaviour, shorter accepted gaps, increased cognitive workload, faster decision cycles, and higher crash risk [30, 32, 11, 39, 40, 35, 36]. Nishant et al. [31] reported that TP significantly increases longitudinal acceleration, braking force variability, and steering fluctuations in a fixed-base simulator, reinforcing that TP elicits statistically significant shifts toward risk-oriented control behaviour.
TP is fundamentally a perception of urgency or limited time to complete the driving task. This perceived urgency narrows the attentional bandwidth and accelerates cognitive processing, often forcing drivers and riders to make rapid decisions with reduced situational awareness [11, 39, 31]. Although early work primarily focused on four-wheelers [28, 11], recent PTW research reveals similar but more sensitive behavioural deterioration under TP, including inconsistent braking, excessive speeding, unstable motor control, and delayed hazard response [39, 32]. These PTW-specific vulnerabilities are important because PTWs exhibit inherently higher crash risk and lower protection levels. Simulator-based studies further show that TP increases crash likelihood across multiple riding tasks. In complex maneuvers such as intersections, Pawar & Velaga [32] reported that TP increases cognitive urgency, speed and reduces caution. By narrowing attention and biasing riders toward fast, risk-oriented strategies, TP directly reduces safety margins and increases the probability of unsafe decisions.
@@ II-B Time Pressure and Gap-Acceptance Behaviour
Research on un-signalized intersections, particularly in LMIC, mainly focused on gap-acceptance behavior and critical gap estimation [19, 27, 32]. Gap acceptance is highly sensitive to TP [32]. Pawar et al. [32] showed that drivers under LTP and HTP accept significantly smaller gaps, resulting in elevated crash likelihoods of 127% (LTP) and 181% (HTP) relative to NTP. Their Generalized Linear Mixed Models (GLMM) and decision-tree (DT) analyses revealed that the TP effects on gap selection are jointly shaped by merge distance, oncoming vehicle speed, speed reduction before merging, and driver experience. These results illustrate how TP systematically modifies perceptual judgment and risk thresholds, thereby reshaping tactical decision-making at intersections.
@@ II-C Time Pressure Effects on PTW Riding: Speed and Over-Speeding
PTWs are particularly vulnerable in traffic, with over-speeding recognized as a major cause of PTW-related crashes [39, 10, 32]. Steg and van Brussel [41] found that subjective norms and favourable attitudes toward speeding significantly increased young moped riders to over-speed, illustrating the role of cognitive and social influences.
Other recent PTW-specific studies show that TP amplifies these vulnerabilities. TP increases over-speeding tendencies due to psychological constructs such as travel-time anxiety, hurriedness, and perceived pressure to reach early [38, 39, 10]. Latent class analyses identify three psychologically distinct rider segments minimally anxious, moderately anxious, and highly anxious, each showing different sensitivity to TP, with highly anxious riders most likely to over-speed. These findings highlight the need to integrate psychological and behavioural information when modelling TP-related riding risk.
@@ II-D Time Pressure and Distraction in MTW Riders
TP represents a cognitive state in which the perceived time available is insufficient to complete the riding task [10, 32]. Under TP, riders must process information more rapidly, often with reduced cognitive reserves. This accelerated decision environment narrows attention, degrades situational awareness, and increases susceptibility to operational errors [29]. Although distraction and TP are distinct constructs, their interaction helps reveal how cognitive stress reshapes rider attention. Gupta et al. [10] found that TP can suppress intentional distractions (e.g., voluntary mobile use) due to heightened task focus, while simultaneously amplifying unintentional cognitive lapses driven by stress and mental workload. These findings reinforce that TP fundamentally alters cognitive processing, attention allocation, and behavioural stability factors central to TP prediction.
@@ II-E Simulator Validity Under Time Pressure
Driving simulators provide a safe, controlled environment for studying TP effects, eliminating real-world collision risks while maintaining experimental validity [31, 7]. Research consistently demonstrates that simulator behavior under TP closely mirrors real-world responses exhibiting what is termed relative validity [31]. These studies show that although absolute behavioral magnitudes may differ between simulated and natural environments, the fundamental patterns of TP-induced changes remain consistent: increased speeds, elevated braking intensity, and amplified vehicular variability occur in both settings. This evidence confirms that simulators are well-suited for TP research when predictive models account for systematic calibration differences between simulated and real-world data.
@@ II-F Machine Learning for Behaviour and Risk Modeling
In parallel to behavioural research, the ITS community has increasingly adopted ML and DL approaches to model driving risk. Traditional classifiers such as Random Forests (RF), Support Vector Machines (SVM), and adaptive Deep Neural Networks (DNNs) have been used to predict high-risk manoeuvres and collision likelihood, often reporting accuracies in the range of 84–90% [37, 1]. More advanced sequence models, including the TST, Informer architecture [49], have demonstrated a strong capability to handle long time-series dependencies relevant to driving dynamics. Despite their predictive success, these models primarily estimate external outcomes (e.g., collisions, hazardous events) and remain largely insensitive to the internal cognitive precursors such as TP, that trigger risky behaviours. This highlights an important methodological gap: existing ML systems detect risk after behavioural deviations occur rather than predicting the underlying cognitive state that gives rise to those deviations.